The billing department of a national cable service company is conducting a study of how customers pay their monthly cable bills. The cable company accepts payment in person at a local​ office, by​ mail, by credit​ card, or by electronic funds transfer from a bank account. The cable company randomly sampled 400 customers to determine if there is a relationship between the​ customer's age and the payment method used. The accompanying sample results were obtained. Based on the sample​ data, can the cable company conclude that there is a relationship between the age of the customer and the payment method​ used? Conduct the appropriate test at the α=0.01 level of significance.

Payment Method 20-30 31-40 41-50 Over 50
In person 9 12 11 12
By mail 29 66 71 56
By credit card 25 19 5 6
By Funds Transfer 23 34 16 6

A) Calculate the test statistic

B) Calculate the p-value

C) Conclusion
Answer #1

Solution:

Here, we have to use chi square test for independence of two categorical variables.

Null hypothesis: H0: There is no relationship between the age of the customer and the payment method ​used.

Alternative hypothesis: Ha: There is a relationship between the age of the customer and the payment method​ used.

We are given level of significance = α = 0.01

A) Calculate the test statistic

Test statistic formula is given as below:

Chi square = ∑[(O – E)^2/E]

Where, O is observed frequencies and E is expected frequencies.

E = row total * column total / Grand total

We are given

Number of rows = r = 4

Number of columns = c = 4

Degrees of freedom = df = (r – 1)*(c – 1) = 3*3 = 9

α = 0.01

Critical value = 21.66599

(by using Chi square table or excel)

Calculation tables for test statistic are given as below:

Observed Frequencies

Column variable

Row variable

C1

C2

C3

C4

Total

R1

9

12

11

12

44

R2

29

66

71

56

222

R3

25

19

5

6

55

R4

23

34

16

6

79

Total

86

131

103

80

400

Expected Frequencies

Column variable

Row variable

C1

C2

C3

C4

Total

R1

9.46

14.41

11.33

8.8

44

R2

47.73

72.705

57.165

44.4

222

R3

11.825

18.0125

14.1625

11

55

R4

16.985

25.8725

20.3425

15.8

79

Total

86

131

103

80

400

Calculations

(O - E)

-0.46

-2.41

-0.33

3.2

-18.73

-6.705

13.835

11.6

13.175

0.9875

-9.1625

-5

6.015

8.1275

-4.3425

-9.8

(O - E)^2/E

0.022368

0.40306

0.009612

1.163636

7.349946

0.618348

3.348329

3.030631

14.67912

0.054138

5.927725

2.272727

2.130128

2.553145

0.926991

6.078481

Chi square = ∑[(O – E)^2/E] = 50.56839

B) Calculate the p-value

P-value = 0.00

(By using Chi square table or excel)

P-value < α = 0.05

So, we reject the null hypothesis

C) Conclusion

There is sufficient evidence to conclude that there is a relationship between the age of the customer and the payment method​ used.